A PARTS file with Part# as hash key includes records with the following Part# values: 2369,...

  1. A PARTS file with Part# as hash key includes records with the following Part# values: 2369, 3760, 4692, 4871, 5659, 1821, 1074, 7115, 1620, 2428, 3943, 4750, 6975, 4981, 9208. The file uses 8 buckets, numbered 0 to 7. Each bucket is one disk block and holds two records. Load these records into the file in the given order using the hash function h(K)=K mod 8. Calculate the average number of block accesses for a random retrieval on Part#.
Answer #1

Solution

How to fill this table

divide each value by 8 then take out the remainder

2369-1
3760-0
4692-4
4871-7
5659-3
1821-5
1074-2
7115-3
1620-4
2428-4
3943-7
4570-6
6975-7
4981-5
9208-0

0

3760

9208

1

2369

2

1074

3

5659

7115

4

4692

1620

2428(OVERFLOW)

5

1821

4981

6

4750

7

4871

3943

6975(OVERFLOW)

OVERFLOW

4692 1620 2428(OVERFLOW) 1821 4981 4750 4871 3943 6975(OVERFLOW)

it allow only 2 posts
but we have overflow at 4 & 7
Please use chaining to solve this

During calculation of average block reads we have 2 block requiring 2 block, and 13 requiring only 1

so
we have

=13/15\*1+2/15*2
=17/15
=17/15 block accesses

---

all the best

if there is anything, please let me know, love to help

please upvote